If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
Answer by ptaylor(2048)

Steven mixes drinks at the rate of 4 drinks per min
Sue mixes drinks at the rate of 2 drinks per min
Jack mixes drinks at the rate of (4/3) drinks per min
Together they mix drinks at the rate of 4+2+(4/3)=12/3 +6/3+4/3=22/3 drinks per min
Let x=amount of time it takes all three working together to mix 20 drinks
Then all three mixes drinks at the rate of 20/x drinks per min
So, our equation to solve is:
22/3 = 20/x or
22x=60
x=2.73 min---------------amount of time it takes all three working together to mix 20 drinks
CK
4*(2.73)+2*(2.73)+(4/3)(2.73)=20
10.9 +5.5+3.6=20
20=20
Hope this helps---ptaylor